### v1

@@ Abstract
Electronic health records (EHRs) enable powerful clinical prediction models but typically offer only coarse, post hoc explanations that are difficult to use in patient-level decision making. To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III. We map ICD-9 codes to PrimeKGchandak2023primekg, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes. We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models on the resulting supervision corpus. On ten PrimeKG-mapped diseases and small training cohorts (400 and 1,000 cases), our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47. Models transfer zero-shot to the CRADLE cohort, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
@@ Introduction
Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
@@
Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLM can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.zhong2024; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions. However, applying CoT in medicine remains challenging. High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability. Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
@@
In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data. We propose a knowledge graph (KG)-guided chain-of-thought framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations. Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation. Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o. Guided by these paths, we prompt an LLM to generate temporally consistent, KG-anchored CoT explanations of whether a disease will appear at the next visit, and we retain only those reasoning traces whose conclusions match the ground-truth label. We then fine-tune lightweight open-weight models on this KG-anchored CoT supervision corpus so that, at inference time, they take ICD-9 features together with disease-specific KG evidence as input and produce both predictions and clinician-style explanations. This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
@@
Task and Data.
We adopt the visit-level next-phenotype prediction task on MIMIC-III described in Section Results, following the 25 phenotyping labels defined by Harutyunyan et al.harutyunyan2019multitask Each index visit at time $t$ is encoded as a binary ICD-9 vector $x_{t}\in\{0,1\}^{7{,}423}$. Among the 25 candidate phenotypes, we focus on a subset $\mathcal{D}$ of 10 diseases that can be reliably mapped into PrimeKG; these 10 labels are the ones used for KG-guided CoT construction and downstream evaluation. For each disease $d\in\mathcal{D}$, we aim to predict whether $d$ will be present at the next visit t+1t{+}1. All models operate on the full ICD-9 feature space, and KG-based variants additionally consume disease-specific evidence derived from PrimeKG as described below.
@@
To incorporate structured biomedical knowledge, we align ICD-9 concepts from MIMIC-III to PrimeKG nodes through a three-stage mapping procedure. Let $E_{\text{ICD}}$ denote the set of unique ICD-9 descriptions, and let $V_{G}$ be the node set of PrimeKG. For each entity $e_{i}\in E_{\text{ICD}}$, we retrieve a similarity-ranked candidate set $S_{i}=\{s_{1},\dots,s_{K}\}\subset V_{G}$ obtained by encoding $e$ and all KG node labels using a text embedding model and ranking candidates by cosine similarity. The final mapped entity $\hat{e}$ is selected in three stages:
Stage 1 (Exact match). If the ICD-9 text of $e_{i}$ exactly matches a node label $s_{k}$ in $S_{i}$, we record $\tilde{e}=s_{k}$.
@@
$\tilde{e_{i}}=\arg\max_{s_{k}\in S_{k}}\cos(e_{i},s_{k})\quad\text{if}\quad\max_{s_{k}\in S_{i}}\cos(e_{i},s_{k})>\tau.$
@@
$\hat{e}_{i}=\mathrm{LLM}(e_{i},\tilde{e}_{i}\mid I_{\text{select}}).$
where $I_{\text{select}}$ is an instruction prompt that asks the LLM to confirm, correct, or reject the candidate mapping based on clinical and biomedical plausibility. This three-stage procedure produces a final set of 1,513 mapped ICD-9 feature (phenotype) nodes and 10 mapped disease nodes, which serve as anchors for downstream relevance mining and reasoning path construction.
@@
Let $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ denote the mapped ICD-9 feature nodes and $\{\hat{e}_{d}\}_{d\in\mathcal{D}}$ denote the mapped disease nodes in PrimeKG (here n=1,513n{=}1{,}513 and $|\mathcal{D}|{=}10$). Each $\hat{e}_{d}$ corresponds to one of the target diseases, while the $\hat{e}_{i}^{f}$ represent potential risk factors, comorbidities, or intermediate biomedical concepts. Our goal is to derive, for each disease $d\in\mathcal{D}$, a compact set of KG reasoning paths linking a small subset of feature nodes to that disease. For disease-centric node relevance, we present the disease node $\hat{e}_{d}$ and the set of all mapped feature nodes $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ to GPT-4o and identify the most relevant features for predicting the presence of $\hat{e}_{d}$ at t+1t{+}1. This yields a disease-level relevance set:
$\mathcal{R}_{d}=\{\hat{e}_{i,d}^{k}\}_{i=1}^{K}=\mathrm{LLM}(\{\hat{e}_{i}^{f}\}_{i=1}^{n},\hat{e}_{d}\mid I_{\mathrm{node\_select}}),\quad K\leq 8\;(\text{we set }K{=}8).$
Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$. Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
$\tilde{\mathcal{P}}_{i,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{i,d}^{k},\hat{e}_{d},G,L\big),$
$\{P_{i,d}^{k}\}_{k=1}^{K}=\mathrm{LLM}\!\big(\tilde{\mathcal{P}}_{i,d},\hat{e}_{d}\mid I_{\mathrm{prune}}\big),\quad K\leq 5\;\;(\text{we set }K{=}5).$
where $\mathrm{shortest\_path}(\cdot,\cdot;G,L)$ denotes the set of all minimum-length paths between two nodes in the biomedical knowledge graph $G$ (PrimeKG), subject to a maximum hop bound $L$ (we use L=5L{=}5 in our experiments) to avoid overly long, weakly supported chains. Since multiple equal-length paths may exist, many of which are clinically irrelevant, weakly supported, or redundant, we apply GPT-4o under a path-selection instruction $I_{\mathrm{path\_select}}$ to obtain a disease-level path set. Aggregating pruned paths across all $\hat{e}_{i}^{f}\in\mathcal{R}_{d}$ yields
$\mathcal{P}_{d}=\mathrm{LLM}\!\big(\{P_{i,d}^{k}\},\hat{e}_{d}\,\big|\,I_{\mathrm{path\_select}}\big),$
a compact subgraph capturing mechanistic or epidemiologic links from relevant clinical concepts to disease $d$. These subgraphs serve as the KG-grounded evidence used in our subsequent CoT generation.
@@
$\mathcal{R}_{d}^{+}(t)$ $=\big\{\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}\;\big|\;x_{t}(\hat{e}_{i,d}^{k})=1\big\},\quad\mathcal{R}_{d}^{-}(t)=\mathcal{R}_{d}\setminus\mathcal{R}_{d}^{+}(t).$
Here $I_{\mathrm{gen}}$ is the CoT instruction prompt, and $x_{t}\in\{0,1\}^{D}$ is the binary ICD–9 feature vector for the index visit at time $t$, with $x_{t}^{+}=\{c\mid x_{t}(c)=1\}$ denoting the set of ICD–9 codes present at that visit. The term $y_{t+1}^{(d)}\in\{0,1\}$ is the ground-truth label indicating whether disease $d$ is present at the next visit t+1t{+}1. The sets $\mathcal{R}_{d}^{+}(t)$ and $\mathcal{R}_{d}^{-}(t)$ indicate which disease-relevant KG nodes mapped from PrimeKG are expressed and not expressed at time $t$, respectively. The condition $x_{t}(\hat{e}_{i,d}^{k})=1$ means that the ICD–9 code corresponding to KG node $\hat{e}_{i,d}^{k}$ is present in the feature vector $x_{t}$.
@@
Each generated chain-of-thought $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$. We use this conclusion to decide whether to keep the example: $C_{t,d}$ is retained if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$. The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs constitutes a compact but high-quality KG-grounded supervision corpus. We fine-tune lightweight open-weight LLMs (e.g., LLaMA, Gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
@@ Results
Dataset:
We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the phenotyping prediction setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits. For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1. This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{D}$ with D=7,423D=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit. The original task is formulated as a multi-label phenotyping problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 phenotyping labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1). For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1. Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict. We perform visit-level splitting over the 12,353 labeled index visits, randomly reserving 10% of cases as a held-out test set. From the remaining pool, we construct training sets of size 400 and 1,000 index visits to study data efficiency, using the rest for model selection.
@@
Evaluation Metrics: Since both the MIMIC-III and CRADLE cohorts exhibit substantial class imbalance, we therefore report accuracy, area under the receiver operating characteristic curve (AUROC), area under the precision–recall curve (AUPR), and macro-F1 as our primary evaluation metrics. For accuracy and F1, we convert predicted probabilities to binary decisions using a fixed threshold of 0.5. Table 1 lists the 10 diseases used as prediction targets.
@@
Table 2: Data-efficiency comparison on MIMIC-III visit-level disease prediction with 400 and 1000 labeled cases (10 diseases). Best metric in bold, second-best underlined.
@@
As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting. With only 400 labeled cases, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$). Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant. When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming strong classical baselines such as XGBoost and Random Forest. These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
@@
In Table 3, MIMIC-trained models are effectively transferred to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes. For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while maintaining similar AUROC ($\approx 0.52$–0.53) and AUPR ($\approx 0.25$). Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to roughly 0.72–0.77 and macro-F1 from 0.31 to around 0.48, again with only modest changes in AUROC and AUPR. Results suggest that the decision boundaries learned from the KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
Human Evaluation:
To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline. For each of 100 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace—one from the KG-guided model and one from the baseline. Without knowing which model produced which trace, clinicians performed pairwise preferences, selecting the better explanation (A or B) along three dimensions:
  • Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions.
  • Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoid unnecessary detail.
  • Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better. The annotation was performed by two human experts from different medical schools: an MD professor and a PhD student in biomedical informatics. In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the PhD student reviewing 55 cases.

### v2

@@ Abstract
Electronic health records (EHRs) enable powerful clinical prediction models but typically offer only coarse, post hoc explanations that are difficult to use in patient-level decision making. To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III. We map ICD-9 codes to PrimeKG, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes. We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models under two resulting training cohorts (400 and 1,000 index visits from MIMIC-III). On ten PrimeKG-mapped diseases and small training cohorts, our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47. Models transfer zero-shot to the CRADLE cohort without any additional fine-tuning, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
@@ Introduction
Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation; zheng2014automatically As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.choi2016retain; ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
@@
Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.wei2022chain; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions. However, applying CoT in medicine remains challenging. High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.xu2024ram; liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability. Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
@@
In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data. We propose a KG-guided CoT framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations. Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation. Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o. Guided by these paths, we prompt an LLM to generate temporally consistent, KG-anchored CoT explanations of whether a disease will appear at the next visit, and we retain only those reasoning traces whose conclusions match the ground-truth label. We then fine-tune lightweight open-weight models on this KG-anchored CoT supervision corpus so that, at inference time, they take ICD-9 features together with disease-specific KG evidence as input and produce both predictions and clinician-style explanations. This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
@@
Prediction Task.
We consider a next visit disease prediction task over structured EHR data. For each patient, let $\{x_{t}\}_{t=1}^{T}$ denotes a sequence of visits ordered in time, where each visit is represented as a binary feature vector $x_{t}\in\{0,1\}^{N}$, with $N$ the dimensionality of the code space. Let $\mathcal{D}$ denotes a set of disease labels of interest, and for each $d\in\mathcal{D}$ and visit index $t$, define a binary outcome $y_{t+1}^{(d)}\in\{0,1\}$, indicating whether disease $d$ is present at the next visit t+1t{+}1. The model’s goal is, for each pair $(x_{t},d)$, to estimate a probability $\hat{p}^{(d)}_{t+1}=f(x_{t},d)$ and an associated binary decision $\hat{y}^{(d)}_{t+1}$ about whether $d$ will appear at visit t+1t{+}1. Our KG-guided models are additionally conditioned on disease-specific KG evidence (relevance sets and reasoning paths) associated with $d$, as described below. In our experiments, we instantiate this general setup using ICD–9 diagnosis codes as features, with N=7,423N=7{,}423 and $|\mathcal{D}|=10$ diseases on the MIMIC-III cohort (details in the Results section).
@@
To incorporate structured biomedical knowledge, we align ICD-9 concepts from MIMIC-III to PrimeKG nodes through a three-stage mapping procedure. Let $E_{\text{ICD}}$ denote the set of unique ICD-9 descriptions, and let $V_{G}$ be the node set of PrimeKG. For each entity $e_{i}\in E_{\text{ICD}}$, we retrieve a similarity-ranked candidate set $S_{i}=\{s_{1},\dots,s_{C}\}\subset V_{G}$ obtained by encoding $e_{i}$ and all KG node labels using a text embedding model and ranking candidates by cosine similarity. The final mapped entity $\hat{e}_{i}$ is selected in three stages:
Stage 1 (Exact match). If the ICD-9 text of $e_{i}$ exactly matches a node label $s_{c}$ in $S_{i}$, we record $\tilde{e}_{i}=s_{c}$.
@@
$\tilde{e}_{i}=\arg\max_{s_{c}\in S_{i}}\cos(e_{i},s_{c})\quad\text{if}\quad\max_{s_{c}\in S_{i}}\cos(e_{i},s_{c})>\tau.$
@@
$\hat{e}_{i}=\mathrm{LLM}(e_{i},\tilde{e}_{i}\mid I_{\text{select}}),$
where $I_{\text{select}}$ is the prompt that asks the LLM to confirm, correct, or reject the candidate mapping based on clinical and biomedical plausibility. This three-stage procedure produces a final set of 1,513 mapped ICD-9 feature nodes and 10 mapped disease nodes, which serve as anchors for downstream relevance mining and reasoning path construction.
@@
Let $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ denote the mapped ICD-9 feature nodes and $\{\hat{e}_{d}\}_{d\in\mathcal{D}}$ denote the mapped disease nodes in PrimeKG (here n=1,513n{=}1{,}513 and $|\mathcal{D}|{=}10$). Each $\hat{e}_{d}$ corresponds to one of the target diseases, while the $\hat{e}_{i}^{f}$ represent potential risk factors, comorbidities, or intermediate biomedical concepts. Our goal is to derive, for each disease $d\in\mathcal{D}$, a compact set of KG reasoning paths linking a small subset of feature nodes to that disease. For disease-centric node relevance, we present the disease node $\hat{e}_{d}$ and the set of all mapped feature nodes $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ to GPT-4o and identify the most relevant features for predicting the presence of $\hat{e}_{d}$ at t+1t{+}1. This yields relevance set:
$\mathcal{R}_{d}=\{\hat{e}_{k,d}\}_{k=1}^{K_{\text{node}}}=\mathrm{LLM}(\{\hat{e}_{i}^{f}\}_{i=1}^{n},\hat{e}_{d}\mid I_{\mathrm{node\_select}}),\quad K_{\text{node}}=8\;,$
where each $\hat{e}_{k,d}$ is a mapped feature node selected from $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ for disease $d$. Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{k,d}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$. Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
$\tilde{\mathcal{P}}_{k,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{k,d},\hat{e}_{d},G,L\big),$
$\mathcal{P}_{d}=\{P_{k,d}\}_{k=1}^{K_{\text{path}}}=\mathrm{LLM}\!\left(\big\{\tilde{\mathcal{P}}_{k,d}:\hat{e}_{k,d}\in\mathcal{R}_{d}\big\},\hat{e}_{d}\,\middle|\,I_{\mathrm{path\_select}}\right),\quad K_{\text{path}}=5,$
where $\mathrm{shortest\_path}(\cdot,\cdot;G,L)$ denotes the set of all minimum-length paths between two nodes in the biomedical knowledge graph $G$ (PrimeKG), subject to a maximum hop bound $L$ (we use L=5L{=}5 in our experiments) to avoid overly long, weakly supported chains. Since multiple paths may exist for a disease, many of which are clinically irrelevant, weakly supported, or redundant, we apply GPT-4o under a path-selection instruction $I_{\mathrm{path\_select}}$ to obtain a disease-level path set, yielding a compact subgraph capturing mechanistic or epidemiologic links from relevant clinical concepts to disease $d$. These subgraphs serve as the KG-grounded evidence used in our subsequent CoT generation.
@@
$\mathcal{R}_{d}^{+}(t)$ $=\big\{\hat{e}_{k,d}\in\mathcal{R}_{d}\;\big|\;x_{t}(\hat{e}_{k,d})=1\big\},\quad\mathcal{R}_{d}^{-}(t)=\mathcal{R}_{d}\setminus\mathcal{R}_{d}^{+}(t).$
Here $I_{\mathrm{gen}}$ is the CoT instruction prompt, and $x_{t}\in\{0,1\}^{N}$ is the binary ICD–9 feature vector for the index visit at time $t$, with $x_{t}^{+}=\{c\mid x_{t}(c)=1\}$ denoting the set of ICD–9 codes present at that visit. The term $y_{t+1}^{(d)}\in\{0,1\}$ is the ground-truth label indicating whether disease $d$ is present at the next visit t+1t{+}1. The sets $\mathcal{R}_{d}^{+}(t)$ and $\mathcal{R}_{d}^{-}(t)$ indicate which disease-relevant KG nodes mapped from PrimeKG are expressed and not expressed at time $t$, respectively. The condition $x_{t}(\hat{e}_{k,d})=1$ means that the ICD–9 code corresponding to KG node $\hat{e}_{k,d}$ is present in the feature vector $x_{t}$.
@@
Each generated CoT $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$. We use this conclusion to decide whether to keep the example: $C_{t,d}$ is kept if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$. The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs forms a compact but high-quality KG-grounded supervision corpus. We fine-tune lightweight open-weight LLMs (e.g., LLaMA grattafiori2024llama, Gemma team2024gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
@@ Results
Dataset.
We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits. For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1. This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{N}$ with N=7,423N=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit. The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 disease labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1). For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1. Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict. We perform visit-level splitting over the 12,353 labeled index visits, randomly reserving 10% of cases as a held-out test set. From the remaining pool, we construct training sets of size 400 and 1,000 index visits to study data efficiency, using the rest for model selection.
@@
Evaluation Metrics. Since both the MIMIC-III and CRADLE cohorts exhibit substantial class imbalance, we therefore report accuracy, area under the receiver operating characteristic curve (AUROC), area under the precision–recall curve (AUPR), and macro-F1 as our primary evaluation metrics. For accuracy and F1, we convert predicted probabilities to binary decisions using a fixed threshold of 0.5. Table 1 lists the 10 diseases used as prediction targets.
@@
Table 2: Data-efficiency comparison on MIMIC-III visit-level disease prediction with 400 and 1000 index visits (10 diseases). Best metric in bold, second-best underlined.
@@
As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting. With only 400 labeled index visits data, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$). Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant. When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming strong classical baselines such as XGBoost and Random Forest. These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
@@
MIMIC-trained models transfer effectively to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes. For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while modestly raising AUROC (to $\approx 0.52$–0.53) and AUPR (to $\approx 0.25$). Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to 0.72–0.77 and macro-F1 from 0.31 to about 0.48, again with modest changes in AUROC and AUPR. These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
Human Evaluation.
To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline. For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace, one from the KG-guided model and one from the baseline. Without knowing which model produced which trace, clinicians made pairwise preferences, selecting the better explanation (A or B) along three dimensions: 1. Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions. 2. Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoids unnecessary detail. 3. Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better. The annotation was performed by two human experts from different medical schools: an MD professor and a MD/PhD student in biomedical informatics. In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the MD/PhD student reviewing 50 cases.